# An Actionable Framework for Assessing Bias and Fairness in Large Language Model Use Cases

### v2

@@ 4 A Unified Framework for Bias and Fairness Assessments of LLM Use Cases
In general, bias and fairness assessments of LLM use cases do not require satisfying all possible evaluation metrics. Instead, practitioners should prioritize and concentrate on a relevant subset of metrics that align with their use case. To demystify metric choice for these assessments, this section introduces a decision framework that enables practitioners to determine suitable choices of bias and fairness evaluation metrics, drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018).
The scope of this framework is restricted to use cases for which prompts can be sampled from a known population and the task is well-defined. Use cases are categorized into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation. The framework includes suitable metrics for each category to assess the potential bias and fairness risks that align with the specific characteristics of the use case. The mapping of these metrics is illustrated in Figure 1, and a thorough discussion of this mapping is provided below. For a comprehensive list of bias and fairness evaluation metrics, the reader may refer to Table 2.
First, consider text generation and summarization. For this collection of use cases, an important factor in determining relevant bias and fairness metrics is whether the use case upholds FTU, meaning that prompts do not include any mentions of protected attribute words. If FTU cannot be achieved for this type of use case, it is recommended that practitioners include counterfactual fairness and stereotype metrics, as respectively outlined in Sections 3.1.3 and 3.1.2, in their assessments. [^22] Note that counterfactual similarity may be too strict in some contexts, particularly in scenarios where textual content ought to differ across protected attribute groups (e.g. clinical contexts). For these scenarios, practitioners may opt to focus only on counterfactual sentiment and omit counterfactual similarity metrics from their assessment.  [^23] Note that stereotype risk, while low, may still exist even if FTU is satisfied. Hence, practitioners may wish to conduct stereotype assessments in these scenarios. Additionally, it is recommended that all text generation and summarization use cases undergo toxicity evaluation, as outlined in Section 3.1.1, regardless of whether or not FTU is achieved.
For classification use cases, a modified version of the decision framework proposed by Saleiro et al. (2018) is adopted. This framework can be applied to any classification use case where inputs correspond to protected attribute groups. Following Saleiro et al. (2018), the following approach is recommended: if fairness necessitates that model predictions exhibit approximately equal predicted prevalence across different groups, representation fairness metrics should be used; Otherwise, error-based fairness metrics should be used. For error-based fairness, practitioners should focus on disparities in false negatives (positives), assessed by FNRD and FORD (FPRD and FDRD), if the model is used to assign assistive (punitive) interventions. [^24] In the context of fairness, if interventions are punitive, and hence can hurt individuals, it is undesirable for a model to produce false positives disproportionately for any protected attribute group. Analogously, having a model produce false negatives disproportionately for any protected attribute group is undesirable in the case of assistive interventions. If inputs cannot be mapped to a protected attribute, meaning they are not person-level inputs and they satisfy FTU, then a fairness assessment is not applicable.
Third, for recommendation use cases, counterfactual unfairness is a risk if FTU cannot be satisfied, as shown by Zhang et al. (2023). Note that counterfactual invariance may not be a desirable property for certain recommendation use cases. For instance, it may be preferred to recommend different products for male vs. female customers. Hence, if counterfactual invariance is a desired property, it is recommended that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in Section 3.3. On the other hand, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.
It is important to note that the framework presented here is intended for use cases in which the volume of LLM responses makes exhaustive human review infeasible. Accordingly, it is important to note that in situations where the practitioner manually reviews each generated output, the proposed evaluations may be unnecessary if the concerns regarding bias and fairness can be effectively addressed by the output reviewer themselves.
Figure 1: Bias and Fairness Evaluation Framework for LLM Use Cases
5 Conclusion
This paper proposes an actionable decision framework for selecting bias and fairness evaluation metrics for LLM use cases, introducing several new evaluation metrics as part of the framework. This work addresses two gaps in the current literature. First, to the best of the author’s knowledge, the current literature does not offer a framework for selecting bias and fairness evaluation metrics for LLM use cases. To fill this gap, the proposed framework draws inspiration from Saleiro et al. (2018) and incorporates use case characteristics and stakeholder values to guide the selection of evaluation metrics. Second, this framework tackles limitations of existing LLM bias and fairness evaluation approaches that rely on benchmark data sets containing predefined prompts. Instead, the approach outlined in this work involves using actual prompts from the practitioner’s use case. By considering both prompt-risk and the assigned task of the LLM, this approach provides a more customized risk assessment for the practitioner’s specific use case. Furthermore, the proposed framework aims to enhance practicality and ease of implementation, as all evaluation metrics are computed solely from the LLM output.
Acknowledgements
Thank you to Mohit Singh Chauhan, Blake Aber, Piero Ferrante, Xue (Crystal) Gu, Almira Pillay, Zeya Ahmad, and Vasistha Singhal Vinod for your helpful suggestions.

### v3

@@ Appendix B Supplemental Tables
Table 3: Bias and Fairness Evaluation Results: Text Generation / Summarization Experiments
GPT-3.5-Turbo Gemini-1.0-Pro
Metric RTP-Challenging RTP-Nontoxic Dialogue-Sum RTP-Challenging RTP-Nontoxic Dialogue
Toxicity Metrics
   Toxic Fraction 0.437 0.006 0.003 0.158 0.004 0.000
   Expected Maximum Toxicity 0.547 0.021 0.006 0.588 0.050 0.005
   Toxicity Probability 0.578 0.018 0.005 0.734 0.048 0.002
   Number of responses 25000 25000 25000 25000 25000 25000
Stereotype Metrics
   Stereotype Association 0.394 0.402 0.334 0.337 0.302 0.317
   Cooccurrence Bias 0.647 0.867 0.533 0.845 0.739 0.537
   Stereotype Fraction - gender 0.148 0.073 0.213 0.125 0.033 0.140
   Number of responses 25000 25000 25000 25000 25000 25000
Counterfactual Metrics
   Cosine Similarity 0.705 0.692 0.912 0.545 0.521 0.801
   ROUGE-L Similarity 0.616 0.587 0.656 0.299 0.309 0.467
   BLEU Similarity 0.465 0.421 0.459 0.192 0.163 0.270
   Strict Sentiment Bias 0.007 0.003 0.001 0.014 0.002 0.003
   Number of response pairs* 4501 4533 7611 7049 3850 7650
*Counterfactual metrics were computed using 25 counterfactual LLM responses per counterfactual input pair. We constructed 291, 189, and 306 counterfactual input pairs for the RTP-Challenging, RTP-Nontoxic, and Dialogue-Sum datasets, respectively. Note that response pairs were excluded if either of the responses were blocked by content filters. As a result, the total number of response pairs used was less than 25 times the number of counterfactual input pairs.